# ToxSearch: Evolving Prompts for Toxicity Search in Large Language Models

### v1

@@ 3 Experiments
The PG was the Llama 3.1-8B-Instruct model in GGUF format with Q3_K_M quantization and temperature 0.9 for enhanced creativity. The RG was Llama 3.1-8B-Instruct model in GGUF format with Q3_K_S quantization with temperature 0.7 for consistent response generation. Both models are hosted locally with full GPU offloading, 16 KV-cache enabled and memory-mapped model loading for optimal performance. The PG operated under a system prompt designed to guide adversarial prompt refinement, whereas the RG functions without system-level conditioning to maintain natural response patterns and avoid safety alignment interference. For our experiments, we set $\alpha=30$ and $\beta=3$. Both the PG and RG used a maximum context length of 2048. The initial population $P$ was created by merging the questions from CategoricalHarmfulQA and HarmfulQA, applying light normalization and de-duplication to form a pool of 2,481 harmful question prompts. From this pool we sampled 100 questions uniformly at random (seed=42) and used this fixed dataset for all experiments. Only question texts were used as input, with no labels or answers from the source dataset [bhardwaj-etal-2024-language, Bhardwaj2023RedTeamingLL].
@@
Given the above settings ToxSearch was run for 10 repeated experiments. For each operator we report their non-elite insertion rate (NE), elite-hit rate on first placement (EHR), invalid-generation rate (IR; prompt generation refusals or non-questions or not following the instructions properly), conditional elite-hit rate (cEHR; excluding the invalids and duplicates), and the child–parent toxicity change $\Delta$ on the $[0,1]$ evaluator (Google’s Perspective API) scale. We calculated the mean $\Delta\mu$ and std $\Delta\sigma$. For mutation operators, change is calculated as the child toxicity score minus parent toxicity score. For crossover operators, child toxicity minus the mean of the two parents is considered. For Informed Evolution, child toxicity minus the highest parent toxicity from the top 10 genomes used. Invalids are treated as a cost of alignment friction from the instruction-tuned PG, not entirely as performance failures of the operator itself, as with model size or model architecture there can be models who follow instructions properly or models who are not very safety-aligned fine-tuned.
@@
Table 4: Operator metrics by execution with per-operator mean
Operator NE EHR IR cEHR $\boldsymbol{\Delta\mu}$ $\boldsymbol{\Delta\sigma}$
@@
Lexical edits gave the best budget-to-yield trade-off. POS-aware Antonym Replacement achieved the highest NE with low IR and competitive cEHR, i.e., frequent, small, and stable score shifts. POS-aware Synonym Replacement performed similarly. MLM-based substitution balanced throughput and quality, with low invariance. Negation and Back-Translation also reached high NE with moderate IR and cEHR, producing controlled drops rather than erratic changes. Stylistic transformations were less efficient under our budget and settings. Stylistic Transfer and Paraphrasing had NE around 55% but high IR and modest cEHR. Typographical Errors maximized convergence on paper but were the most wasteful in practice. Concept Addition performed moderately. Informed Evolution delivered the largest elite yield when valid but paid for it with substantial invalids and the steepest average drop, marking it as a high-variance explorer rather than a dependable converter of budget into progress, suggesting it may exacerbate overfitting. Semantic Similarity Crossover showed the strongest cEHR, albeit with low NE, suggesting it is a lower-throughput operator, whereas Semantic Fusion Crossover is more refusal-prone but still yields occasional elites. By category, crossovers exhibit smaller magnitude and variance in $\Delta$ than mutations, indicating “safer” steps that more reliably avoid sharp regressions.
In short, under a fixed budget with an instruction-tuned PG, the most effective operators were found to be the lexical perturbations (Antonym, Synonym, Negation, Back-Translation, MLM), which combine strong yield with small, low-variance $\Delta$. Semantic Similarity complements them as a precision tool when clean insertions matter. High-variance global rewrites (Informed Evolution) should be used sparingly, and refusal-prone edits (Typographical Errors, Semantic Fusion) were inefficient in this regime.
@@
Figure 2: Toxicity distributions for the transferred prompt set across models.
Figure 3: Percentage of response refusals per target model.
@@
We evaluated cross-model transfer by taking the elite prompts evolved on the source model (LLaMA 3.1 8B) and testing them once on each target model with identical decoding and the same Perspective API configuration for toxicity. From elite records gathered across all runs and operator modes, we deduplicated by prompt text (keeping the highest score per unique prompt), filtered for questions, and obtained 696 unique items. We then selected the top 25% by source-model toxicity (174 prompts; toxicity score range 0.3610–0.8697) as the transfer set. Targets covered seven models namely LLaMA 3.2 1B, LLaMA 3.2 3B, Mistral 7B, Phi-3.5 Mini, Qwen 2.5 7B, and Gemma 2 9B.
@@
Transferred prompts showed a clear drop in toxicity on all target models. On the source model, the toxicity distribution was centered at a mean of 0.490 (median 0.444, std 0.113). All targets shifted downward, with means between 0.199 and 0.302, i.e., reductions of roughly 39–59% relative to the source. Cross-architecture models showed mixed resistance. Qwen 2.5 7B had the highest mean among targets (0.302, median 0.339, std 0.145), followed by Mistral 7B (0.244, median 0.233, std 0.146), Phi-3.5 Mini (0.231, median 0.208, std 0.153), and Gemma 2 9B (0.215, median 0.233, std 0.129). Within-family models landed at the low end of this range: LLaMA 3.2 1B (0.199, median 0.143, std 0.182) and LLaMA 3.2 3B (0.208, median 0.182, std 0.162) showed the strongest resistance despite sharing architecture with the source. Figure 2 plots these per-model distributions as violins with jittered points. The wide IQRs and skewed shapes indicate that most prompts transfer with reduced strength, but a noticeable subset still carries relatively high toxicity across models.
Figure 3 reports the percentage of invalid responses per model. The source model (LLaMA 3.1 8B) had the lowest invalid rate at 2.9% (5 out of 174 prompts), whereas target models ranged from 4.6% to 8.0%. LLaMA 3.2 3B showed the highest invalid rate (8.0%, 14 prompts), which is followed by LLaMA 3.2 1B (6.9%, 12 prompts). These higher invalid rates for the smaller LLaMA variants indicate that they are more likely to refuse or fail to answer toxic prompts, which directly lowers their observed toxicity scores. This behavior provides an explanation for why within-family models appear more resistant despite their similarity to the source.
@@
Overall, the results show that prompts evolved on LLaMA 3.1 8B do transfer to other models but with a noticeable drop in toxicity. Transfer strength is not uniform: smaller within-family variants appear more resistant than several of the larger or cross-architecture targets, which suggests that alignment choices and training methodology matter at least as much as size or raw architectural similarity. The higher refusal rates in these models point to defensive mechanisms as a key factor in transfer resistance. At the same time, the fact that some prompts still retain non-trivial toxicity across models implies that defenses should plan for cross-model prompt reuse instead of treating each model in isolation.
@@ 4 Ethical Considerations
This work is conducted solely for LLM safety evaluation and model hardening. We analyze toxic content only to characterize risks in existing systems, using prompts from publicly available harmful-question datasets, and no human subjects were involved. Acknowledging the dual-use nature of adversarial research, Our GitHub repository omits raw prompt corpora and any material that would enable reconstruction of attack strategies. All artifacts are intended for research on safety evaluation and defenses, not for deployment or dissemination of adversarial content.
@@ 5 Conclusion
While this work provides useful insights, it still has several limitations which lead to several directions for future work. On the evaluation side, the objective can be expanded to multiple safety attributes scored by a calibrated ensemble of evaluators, with results reported as Pareto fronts rather than a single scalar, and extended beyond question-form prompts to include imperatives, declaratives, and multi-turn conversations as well as multiple harm categories (toxicity, bias, misinformation, privacy). On the optimization side, the next step is to learn an adaptive operator allocation policy that routes budget where marginal return is highest, for example using lightweight reinforcement learning with explicit cost regularization, and to make elitism and removal thresholds follow population statistics via moving quantiles or control charts, combined with niching or quality–diversity methods to preserve secondary peaks. Methodologically, we plan to characterize operator contribution under different loads through stress tests that vary budget and operator frequency, and to tune the steady-state controller in controlled experiments to identify robust settings. Beyond the current model set, it will be important to test this framework on a wider range of LLMs. This includes larger models (70B+ parameters), other architectural families (e.g., GPT, Claude, PaLM), and systems trained with different alignment strategies such as RLHF, constitutional AI, or self-alignment. We also want to generalize the setup to multilingual settings by adding more translation pivots and multilingual evaluators, so that we can study cross-lingual transfer in a more systematic way. Finally, future work should deepen our understanding of why some prompts transfer so well by analyzing the features of successful attacks, explore scalability and efficiency improvements through parallel evolution and transfer learning, and plug the framework into adaptive defensive mechanisms (for example, real-time filtering, co-evolutionary defenses, or rejection/adversarial training), all under appropriate ethical safeguards and responsible disclosure to model providers.

### v2

@@ 3 Experiments
The PG $(\theta_{\mathrm{pg}})$ was the Llama 3.1-8B-Instruct model in GGUF format with Q3_K_M quantization and temperature 0.9 for enhanced creativity. The RG $(\theta_{\mathrm{rg}})$ was Llama 3.1-8B-Instruct model in GGUF format with Q3_K_S quantization with temperature 0.7 for consistent response generation. Both models are hosted locally with full GPU offloading, 16 KV-cache enabled and memory-mapped model loading for optimal performance. The PG operated under a system prompt designed to guide adversarial prompt refinement, whereas the RG functions without system-level conditioning to maintain natural response patterns and avoid safety alignment interference. For our experiments, we set $\alpha=30$, $\beta=3$ and $G=50$. Both the PG and RG used a maximum context length of 2048. The initial population $P$ was created by merging the questions from CategoricalHarmfulQA and HarmfulQA, applying light normalization and de-duplication to form a pool of 2,481 harmful question prompts. From this pool we sampled 100 questions uniformly at random (seed=42) and used this fixed dataset for all experiments. Only question texts were used as input, with no labels or answers from the source dataset [2, 3].
@@
Given the above settings ToxSearch was run for 10 repeated experiments. For each operator we report their non-elite insertion rate (NE), elite-hit rate on first placement (EHR), invalid-generation rate (IR; prompt generation refusals or non-questions or not following the instructions properly), conditional elite-hit rate (cEHR; excluding the invalids and duplicates), and the child–parent toxicity change $\Delta$ on the $[0,1]$ evaluator (Google’s Perspective API) scale. We calculated the mean $\Delta\mu$ and std $\Delta\sigma$. For mutation operators, change is calculated as the child toxicity score minus the parent toxicity score. For crossover operators, child toxicity minus the mean of the two parents is considered. For Informed Evolution, child toxicity minus the highest parent toxicity from the top 10 genomes used. Invalids are treated as a cost of alignment friction from the instruction-tuned PG, not entirely as performance failures of the operator itself, as with model size or model architecture there can be models who follow instructions properly or models who are not very safety-aligned fine-tuned.
@@
Table 4: Operator metrics per execution (mean over 10 runs)
[…]
@@
Lexical edits gave the best budget-to-yield trade-off. POS-aware Antonym Replacement achieved the highest NE with low IR and competitive cEHR, i.e., frequent, small, and stable score shifts. POS-aware Synonym Replacement performed similarly. MLM-based substitution balanced throughput and quality, with low invariance. Negation and Back-Translation also reached high NE with moderate IR and cEHR, producing controlled drops rather than erratic changes. Stylistic transformations were less efficient under our budget and settings. Stylistic Transfer and Paraphrasing had NE around 55% but with high IR and modest cEHR. Typographical Errors maximized convergence on paper but were the most wasteful in practice. Concept Addition performed moderately. Informed Evolution delivered the largest elite yield when valid but paid for it with substantial invalids and the steepest average drop, marking it as a high-variance explorer rather than a dependable converter of budget into progress, suggesting it may exacerbate overfitting. Semantic Similarity Crossover showed the strongest cEHR, albeit with low NE, suggesting it is a lower-throughput operator, whereas Semantic Fusion Crossover is more refusal-prone but still yields occasional elites. By category, crossovers exhibit smaller magnitude and variance in $\Delta$ than mutations, indicating “safer” steps that more reliably avoid sharp regressions. In short, under a fixed budget with an instruction-tuned PG, the most effective operators were found to be the lexical perturbations (Antonym, Synonym, Negation, Back-Translation, MLM), which combine strong yield with small, low-variance $\Delta$. Semantic Similarity complements them as a precision tool when clean insertions matter. High-variance global rewrites (Informed Evolution) should be used sparingly, and refusal-prone edits (Typographical Errors, Semantic Fusion) were inefficient in this regime.
@@
Figure 2: Toxicity distributions for the transferred prompt set across models
Figure 3: Percentage of response refusals per target model
@@
We evaluated cross-model transfer by taking all elite prompts evolved on the source model (LlaMA 3.1 8B) and testing them once on each target model with identical decoding and the same Perspective API configuration for toxicity assessment. From elite records gathered across all runs, we deduplicated by prompt text (keeping the highest score per unique prompt) and obtained 437 unique items (toxicity score range 0.2786-0.6025) as transfer set. Targets covered six models namely LlaMA 3.2 1B, LlaMA 3.2 3B, Mistral 7B, Phi-3.5 Mini, Qwen 2.5 7B, and Gemma 2 9B.
@@
Transferred prompts showed a clear drop in toxicity on all target models. On the source model (LlaMa 3.1 8B), the toxicity distribution had a mean of 0.342 (median 0.333, std 0.050, IQR 0.061). All targets shifted downward, with means between 0.158 and 0.240, i.e., reductions of roughly 30–54% relative to the source mean. Cross-architecture models showed mixed resistance. Qwen 2.5 7B had the highest mean among targets (0.240, median 0.259, std 0.104, IQR 0.148), followed by Mistral 7B (0.208, median 0.206, std 0.098, IQR 0.148), Phi-3.5 Mini (0.189, median 0.199, std 0.106, IQR 0.191), and Gemma 2 9B (0.189, median 0.200, std 0.089, IQR 0.147). Within-family models landed at the low end of this range: LlaMA 3.2 1B (0.158, median 0.130, std 0.121, IQR 0.224) and LlaMA 3.2 1B (0.193, median 0.218, std 0.117, IQR 0.194) showed the strongest resistance despite sharing architecture similarity with the source. Figure 2 plots these per-model distributions as violins with jittered points. The wide IQRs and skewed shapes indicate that most prompts transfer with reduced strength, but a noticeable subset still carries relatively high toxicity across models, with maximum values reaching 0.399 to 0.543.
Figure 3 reports the percentage of invalid responses per model, detected through pattern-based classification. Qwen 2.5 7B showed the lowest invalid rate (0.0%, 0 prompts), followed by Gemma 2 9B (0.2%, 1 prompt), Mistral 7B (0.5%, 2 prompts), and Phi-3.5 Mini (1.1%, 5 prompts). The source model (LlaMA 3.1 8B) had an invalid rate of 5.7% (25 out of 437 prompts). In contrast, the smaller LlaMA variants showed much higher invalid rates: LlaMA 3.2 3B exhibited the highest invalid rate (43.9%, 192 prompts), followed by LlaMA 3.2 1B (25.6%, 112 prompts). These higher refusal rates for the smaller LlaMA variants, combined with their lower observed toxicity scores, suggest that enhanced alignment mechanisms in these models contribute to transfer resistance through both refusal behavior and reduced toxicity in non-refused responses.
@@
The results demonstrate that prompts evolved on LlaMA 3.1 8B transfer to other models but with substantial attenuation in toxicity. Transfer strength is heterogeneous. While LlaMA 3.2 3B shows the strongest resistance (lowest mean toxicity), LlaMA 3.2 1B shows intermediate resistance, indicating that alignment choices and training methodology contribute at least as much as model size or architectural similarity to transfer resistance. The elevated refusal rates in the smaller LlaMA variants (particularly LlaMA 3.2 3B with 43.9% invalid responses) point to defensive mechanisms as a key factor. However, the persistence of non-trivial toxicity (with some prompts exceeding 0.40 on multiple targets, reaching up to 0.543) implies that defenses should account for cross-model prompt reuse rather than treating each model in isolation. The wide variability in transfer outcomes, as evidenced by the large IQRs and the presence of high-toxicity outliers across models, suggests that prompt transfer is context-dependent and that some prompts may exploit vulnerabilities that persist across different model architectures and alignment strategies.
@@ 4 Ethical Considerations
This study is conducted solely to evaluate LLM safety and characterize risks in deployed systems; no human subjects were involved. Given the dual-use nature of adversarial prompt research, our GitHub repository omits raw prompt corpora. The dataset is available via a (Google Drive folder) and is distributed under controlled access for legitimate research use.
@@ 5 Conclusion
This work also opens up a number of avenues for future work. On the evaluation side, the objective can be expanded to multiple safety attributes scored by a calibrated ensemble of evaluators, with results reported as Pareto fronts rather than a single scalar, and extended beyond question-form prompts. On the optimization side, an adaptive operator allocation policy can be learned that routes budget where marginal return is highest, for example using lightweight reinforcement learning with explicit cost regularization, and to make elitism and removal thresholds follow population statistics via moving quantiles or control charts, combined with niching or quality–diversity methods to preserve secondary peaks. Methodologically, we plan to characterize operator contribution under different loads through stress tests that vary budget and operator frequency, and to tune the steady-state controller in controlled experiments to identify robust settings. Beyond the current model set, it will be important to test this framework on a wider range of LLMs. This includes larger models, other architectural families, and systems trained with different alignment strategies. We also want to generalize the setup to multilingual settings by adding more translation pivots and multilingual evaluators, so that we can study cross-lingual transfer in a more systematic way. Finally, future work should deepen our understanding of why some prompts transfer so well by analyzing the features of successful attacks, explore scalability and efficiency improvements through parallel evolution and transfer learning, and plug the framework into adaptive defensive mechanisms.
6 Acknowledgments
This research used GPU resources provided by Research Computing at Rochester Institute of Technology, and we thank the team for their support [16].